OpenStudy (anonymous):

Jackie Rich would like to buy a $19,700 Toyota hybrid car in 5 years. Jackie wants to put the money aside now. Jackie’s bank offers 10% interest compounded semiannually. How much must Jackie invest today?

OpenStudy (anonymous):

I don't have the calculator to do the exponents, but I can help you out to that point.

OpenStudy (anonymous):

The equation I looked up said A = P(1 + r/n) ^nt

OpenStudy (anonymous):

A is for ending amount so 19700 P is beginning amount which is unknown r is interest rate (as decimal) so .1 n is number of compounds a year, so 2 t is total number of years \[19700=P(1+\frac{ .1 }{ 2 })^{2(5)}\]

OpenStudy (anonymous):

thank you so much!

OpenStudy (anonymous):

Your welcome, but just in case it gave me another equation to help with this. Do you think you will need it?

OpenStudy (anonymous):

\[\frac{ A }{ (1+\frac{ r }{ n })^{nt} }=P\]